OpenStudy (anonymous):

3wx/6x multiplied by 3wx/9w express product in simplest form

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

how long is the answer

OpenStudy (skullpatrol):

$$\Huge \frac{3wx}{6x}*\frac{3wx}{9w}$$

OpenStudy (anonymous):

You just restated my question I need the answer

OpenStudy (anonymous):

lol

OpenStudy (skullpatrol):

$$\Huge Answer \neq Understanding$$

OpenStudy (anonymous):

yeah thanks for all the help but still wrong answer

OpenStudy (anonymous):

w^3 x^3/6

OpenStudy (anonymous):



OpenStudy (anonymous):

its either wx/6, wx/3, 2/3(wx), or wx

OpenStudy (anonymous):

those are the only choices I have

OpenStudy (skullpatrol):

$$\Huge \frac{3wx}{6x}*\frac{3wx}{9w}$$

OpenStudy (anonymous):

not sure if its wrong or right cause platos wrong sometimes

OpenStudy (anonymous):

its probably plato

OpenStudy (anonymous):

alex so u going to your next class just curious im not sure if im going or not

OpenStudy (skullpatrol):

$$\Huge \frac{3wx}{6x}*\frac{3wx}{9w}$$ $$\Huge \frac{3w\cancel{x}}{6\cancel{x}}*\frac{3\cancel{w}x}{9\cancel{w}}$$

OpenStudy (anonymous):

wx/6